OpenStudy (anonymous):

Evaluate these integrals using integration by parts? a) ∫ cos (In x) dx b) ∫ x tan^(2) x dx

OpenStudy (atlchic):

what u say first

OpenStudy (anonymous):

Please? :)

OpenStudy (atlchic):

tell me guess

OpenStudy (anonymous):

I'm not sure how to do these questions :(

OpenStudy (watchmath):

a) Let \(u=\cos(\ln x)\) and \(dv=dx\). So \(du=-\frac{1}{x}\sin(\ln x)dx\) and \(v=x\). So the integral is equal to \(uv-\int vdu\) \(x\cos(\ln x)+\int\sin (\ln x)\)dx Do integration by parts one more time to get \(\int\cos(\ln x)\,dx=x\cos(\ln x)+x\sin(\ln x)-\int \cos(\ln x)\,dx\) Then we have \(2\int\cos(\ln x)\,dx=x\cos(\ln x)+x\sin(\ln x)\) Therefore \(\int\cos(\ln x)=\frac{x}{2}(\cos (\ln x)+\sin (\ln x))\)+C.

OpenStudy (anonymous):

second one i think is the same. \[u=x\] \[du=dx\] \[dv = tan^2(x)\] \[v=tan(x) -x\] \[\int x tan^2x dx=x(tan(x)-x)-\int (tan(x)-x)dx\] \[=xtan(x)-x^2+ln(cos(x))+\frac{x^2}{2}\]